OpenStudy (anonymous):

(1-sin(pi/8))(1+sin(pi/8))

OpenStudy (anonymous):

1-sin^2(pi/8)

OpenStudy (anonymous):

1+sin(pi/8)-sin(pi/8)-sin^2(pi/8)

OpenStudy (anonymous):

=cos^2(pi/8)

OpenStudy (anonymous):

And that is equal? How to find out :) ?

OpenStudy (whpalmer4):

\[\cos^2x+\sin^2x = 1\]\[1-\sin^2x = \cos^2x\]

OpenStudy (whpalmer4):

You want to evaluate \(\cos^2(\pi/8)\)?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

How to find out what cos2(π/8) is

OpenStudy (anonymous):

cos^2(pi/8)

OpenStudy (whpalmer4):

Okay, that's easy enough. Use the identity \[\cos^2 u = \frac{1+\cos 2u}{2}\]

OpenStudy (whpalmer4):

\(u = \pi/8\) so \(2u = \pi/4\) and you can find the \(\cos\) of that easily.

OpenStudy (anonymous):

cos^2x=(1+sqr2/2)/2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

=(2+sqr2)/4

OpenStudy (whpalmer4):

That's correct. About 0.853553 if evaluated.

OpenStudy (anonymous):

Tyvm :)

OpenStudy (whpalmer4):

yw :)

OpenStudy (whpalmer4):

study those identity lists!

OpenStudy (anonymous):

Will do :D